Julie worked 32 hours last week and was paid $256. At the same rate,
what will Julie be paid if she works 40 hours this week?

Answer by stanbon(75887)

Julie worked 32 hours last week and was paid $256. At the same rate,
what will Julie be paid if she works 40 hours this week?
-----
Solve:
x/40 = 256/32
----
x = 40(256/32)
---
x = (5/4)256
----
x = 5*64
----
x = $320
==========
Cheers,
Stan H.
==========